OpenStudy (anonymous):

find the distance between 2,0 and -2,8

OpenStudy (anonymous):

do the distance formula \[\sqrt(x2 −x1 )^ 2 + (y2 - y1 )^2 \]

OpenStudy (anonymous):

i get a decimal tho

OpenStudy (anonymous):

then plug in the points (2,0) and (-2,8) it will be\[\sqrt(-2 - 2)^2 + (8 - 0)^2 =\sqrt(-4)^2+(8)^2 \] \[\sqrt( 16) + (64) =\sqrt (80)\]

OpenStudy (anonymous):

so the simplified version of \[\sqrt(80) \] is \[4\sqrt(5)\]

OpenStudy (anonymous):

thats the answer right?

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

if you need help ask me

OpenStudy (anonymous):

find the intercepts and use them to sketch the graph... 3y-5x=0

OpenStudy (anonymous):

well the y intercept is going to be 3y = 5x then y = 5/3x and x is going to be -5x = -3y then x = -3/-5y which will be x = 3/5y

OpenStudy (anonymous):

and plug them in and you get the graph

OpenStudy (anonymous):

wait no!!!!!! I was just blanked out over there when i wrote.

OpenStudy (anonymous):

ok wait almost got it

OpenStudy (anonymous):

Okay, you need to find the intercepts, correct? Here's how I do it: Set x = 0, solve for y: 3y−5(0)=0 3y=0 y= Set y = 0, solve for x: 3(0)−5x=0 −5x=0 x= The point where x = 0 is the y-intercept. The point where y = 0 is the x-intercept.

OpenStudy (anonymous):

so the intercepts are (0,0)